Res) flant Gain / 425t 2 Ca) wlen PID-Contrel Gainsks) Klsts)(it Spetel the Yoot - locus (Plot the keoKeoe) : abnt oPen-loop Plart's respause YS)-GCS)us) /(mit step Tu) x ax3:time t) Plot the g raph (usng muene (aplace fam) Y RTS pesfonse YIE) ConPare on a bant Closes-loop system's resparse ys= T19)r)/(gan k = 14) Plot the Yaxis: pespaye E aPh Cushy inere larhce torm) L4:CGiven Gis)SH Plot the root 10cus on whan p-lantnGain is es)=k Pectively aud Po-Cantrol Gain is ks)=K (S+2)
Res) flaunt Cain s42512 La) When PID-Contrel Gain s Ks) Klsts)(it) Gpoteh the Yoot - acus (Plot the KeoKeco) : abnt ofen-10p Plant's respanse (s) -GCS)us)/mit step Tp) x aT3:Ene r) Plot the graph (usng invene (aplace fam) Y ATS resfonse Ye) Cu Pare on abnt Closel -loop system's resparse 1S=TIS)rs)/(Jank=14) Plot the yaris pesae yePh luhg mere larhce torm) L): Caiven Gis) Plot tha rot lous whan P-lntnGain is Kcs)-k Jrapeiey aud PD-antrol Gain is ks)=K (st2)
Answer #1

first page of solution is uploaded at last.uploaded pages of order is 2,3,4,5,6,7,8,9,1. The page number is written at the top right corner of the each page.please adjust.if you have any doubt comment to me.thank you.

(-1+15) C1t1t0-2) (2 26H S-1t) t an tan +tan -90 14.036-51.34-9 +4s 2.304 lab7 ( 80 82.3oq 97 696 97.696 PLA ko - 0.2 420 - 5as2) Cs4525t1)- s34241s) C2s+52) dind 2 0-8s+ 4S +26t lo 4st) 84+ AS.6 S+ ss+ 4s -2s-5. 2s3 453-10-45-4sY-164 sニo S9 10.45 +1ftst) - +5Csti) 1) t aj (P-C22) 25cst1) +2 ytu aj (1-1 Ct25t 2) yesr S41t) + 1 2 -2 (s+2St2 S+2 20 S+1 I 2 S 2 St1) + 1 Mpply14. KY42S42 k. 6cs) Ces) TIS) - KGCS) RCS) 14 425t Ces) TOS): S+2S+16 RCS) 14 4ESE TLS) RCS) RIS) sCs2st16) SC s+25+16) s(S+114 yes) SCチ74)-299538 S1+53.89 14 S+1-f3.89 了チ44+29.9538 +2 9.95 6) + S+1) +J 3 82)(474-299536) 8 S 7 2St1 RC2.9538) Cst)+ 2(tzo -1 o.12S 七こ1 → yu)2 O.367 0.9971-0.0675 - 0.51S 9]= te5 44t)= - 0. 006 0.943-0 085 = 0.869 tz 1o t) 0-7801-0.1(6LS 0-8749GUS Ststi) i + kcs)E k. Open Loop T.F: Sts+1) No.0 poler z, No-o zeros ニ2. No-of tinite No.of RLB poles poles zevos L No.ot AIf kLs) kCst 2) KCst2) Open oop TF = SiStI) No.0 poles = 2 No. of zeTOS RetaB (N) P-2) = , Asymptodes. = 180 9o No of にこD GEGplant Gain a tet 4 assume Negative feedback GeD Gain (s+s)Ct K Cs+5) Cs+0-2) Gain Open loop TF kCs+5)(s+o.2) て+Sてよ KCS+5) CSt